OpenStudy (anonymous):

nine times the sum of 4 and some number is 198. what is the number?

OpenStudy (anonymous):

CR7 rocks

OpenStudy (anonymous):

how did u get that?

OpenStudy (anonymous):

9(4+x)=198

OpenStudy (anonymous):

thx i know my favorite player

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

Me too a big fan of him

OpenStudy (anonymous):

do u multiply 9 and 4 or divide them into 198

OpenStudy (anonymous):

9(4+x)=198 9*4+9x=198 36+9x=198 9x=198-36 9x=162 x=162/9 x=18

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

oh ok i got it

OpenStudy (anonymous):

Great

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

viva ronaldo